# Learning to Control Unknown Strongly Monotone Games

### v2

@@ 2.1.1 Quadratic Global Objective
For a positive semi-definite matrix $\boldsymbol{G}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{\rho}\in\mathbb{R}^{Nd}$, the controller can set a global cost of the form $\Phi\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$. All the action profiles that are local optima of $\Phi\left(\boldsymbol{x}\right)$ satisfy $\nabla\Phi\left(\boldsymbol{x}\right)=2\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}=0$. Hence, using our approach, we can set $\boldsymbol{A}=2\boldsymbol{G}$ and $\boldsymbol{\ell^{*}}=-\boldsymbol{\rho}$ to reach a global optimum.
@@ 2.1.2 Weighted Resource Allocation Games
In resource allocation, high congestion on a single resource increases the operation costs for the manager, can lead to system failures, and incurs suboptimal performance for the users. Therefore, load balancing is a key control objective for the manager that supervises the resource allocation protocol. In the electricity grid, the manager can be the utility company running Demand Side Management (DSM) to cut the production costs of energy at peak hours. In wireless networks, the manager can be an access point coordinating a protocol that minimizes the power consumption of the devices and the interference in the network. Other examples are load balancing in data-centers [55] and control of parking resources [56]. For distributed resource allocation, designing an efficient load balancing protocol becomes far more challenging.
In weighted resource allocation, each dimension of a player’s action represents how much they use from each of the $d$ resources. We have $d$ linear constraints i.e., $K=d$ and each represents a target load for each resource $i$, i.e.,
@@
The manager can set the target loads $\boldsymbol{l}^{*}$ to optimize application-specific objectives such as system efficiency, regulations, and operation costs. The controlled coefficients $\boldsymbol{\beta}$ represent incentives to use a resource. The weights in $\boldsymbol{A}$ are based on how each agent’s use affects the resource.
A special case is resource allocation without weights, studied in [50]. Generalizing to a weighted sum of actions allows modeling of power control in wireless networks and load balancing in data centers. In wireless networks, the interference that the manager measures is a weighted sum of the transmission powers of the players, where the weights are the channel gains. In data centers, different users send jobs of different sizes to different servers (resources), so the load is a weighted sum of their actions (how many jobs to send where).
@@ 3 Game Control Algorithm
Algorithm 1 Online Game Control with Bandit Feedback
@@
    The manager observes the vector $\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}}$.
@@
The main challenge in guaranteeing that the dynamics converge to a point where the linear constraints are satisfied is that the manager does not know the game, consisting of $\left\{r_{n}\right\}$ and $\left\{\mathcal{X}_{n}\right\}$. To overcome this, we propose an online approach where the manager learns how to adjust $\boldsymbol{\alpha}_{t}$ to satisfy the linear constraints by using the iteration in (8) with the control step-size sequence $\left\{\varepsilon_{t}\right\}$. We only assume that the manager can observe the instantaneous constraint violation $\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}}$ at the start of turn $t$. This feedback maintains basic privacy for the players since the manager does not know their reward functions and does not even observe their individual actions.
Our algorithm is a two time-scale stochastic approximation algorithm. In the faster time-scale, dictated by the stepsize $\{\eta_{t}\}$, the agents update their action using stochastic gradient descent, taking a step towards the NE corresponding to $\boldsymbol{\alpha}_{t}$. In the slower timescale, dictated by the stepsize $\{\epsilon_{t}\}$, the manager observes the constraint violation based on the players’ current action updates $\boldsymbol{\alpha}_{t}$ to steer the player towards a desirable point.
Essentially, our algorithm is an iterative approach to solving the constrained optimization problem using Lagrange multipliers. We wish to solve the optimization problem with gradients given by $F(\boldsymbol{x})$ under the constraints $\boldsymbol{x}\in\mathcal{X}$ and $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$. The agents cannot ensure that $\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}}$ is satisfied for all $t$ without communicating, and hence the manager uses an iterative approach instead. Similar approaches have been used in [57, 58] for different problems.
Our players run stochastic gradient descent to optimize their rewards, and therefore are myopic and do not look to manipulate the manager using dynamic strategies. This behavior can be encountered with selfish players in a large-scale network or with cooperative players.
In a selfish setting, players are likely to maximize their rewards using a widespread online learning algorithm such as gradient descent [59, 10]. In particular, when the players do not know their reward functions, the small steps taken by SGD allow for gradual learning, as opposed to best-response dynamics that can have large jumps. Nevertheless, our techniques can be applied to other distributed algorithms that are guaranteed to converge to NE.
In a cooperative setting, the behavior of the players is not something to model but to design [60]. The algorithm is used as a distributed protocol programmed into the devices in the network (e.g., WiFi). The challenge is then not to control the selfish behavior of the players, but to guide them towards a more efficient NE they do not have enough information to find otherwise (i.e., players do not know which $\boldsymbol{\alpha}$ would satisfy the linear constraints at NE). For this purpose, gradient play is an efficient distributed algorithm, that can lead to an efficient NE when combined with the manager’s algorithm in (8).
@@ 3.1 Performance Guarantees
The manager wants to converge to a suitable $\boldsymbol{\alpha}$ such that the linear constraints in (6) are satisfied at the NE corresponding to $\boldsymbol{\alpha}$, i.e., converge to the set
@@
We make the following assumption on our game to ensure that $\mathcal{N}_{opt}$ is non-empty, which is just Slater’s condition [61] for convex optimization.
@@
The action set is of the form $\mathcal{X}=\{\boldsymbol{x}\mid h_{i}(\boldsymbol{x})\leq 0,i=1,\ldots,M\}$ for some $M$ where each $h_{i}(\boldsymbol{x})$ is continuously differentiable and convex and there exists $\boldsymbol{x}\in\text{int}(\mathcal{X})$ (i.e., $h_{i}(\boldsymbol{x})<0$, $i=1,\ldots,M$) such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$.
@@
The stepsizes $\{\eta_{t}\}$ and $\{\epsilon_{t}\}$ are of the form $\eta_{t}=1/(t+T_{1})^{\eta}$ and $\epsilon_{t}=1/(t+T_{2})^{\epsilon}$ for some $T_{1},T_{2}>0$ which satisfy:
  1. (a)
    $\{\eta_{t}\},\{\epsilon_{t}\}$ are monotonically decreasing sequences that are non-summable, but square summable.
  2. (b)
    $\lim_{t\to\infty}\frac{\epsilon_{t}^{2}}{\eta_{t}^{3}}=0$.
Our first result shows that $\boldsymbol{\alpha}_{t}$ converges to $\mathcal{N}_{opt}$ and hence $\boldsymbol{A}\boldsymbol{x}_{t}$ converges to $\boldsymbol{\ell^{*}}$.
@@
Under Assumptions 1-5, $\boldsymbol{\alpha}_{t}$ converges to the set $\mathcal{N}_{opt}$, $\boldsymbol{x}_{t}$ converges to $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$, and
$\lim_{t\rightarrow\infty}\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}},$
with probability 1.
The following theorem gives finite-time convergence guarantees for our algorithm.
@@
Under Assumptions 1-5,
$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(\eta_{t}+\frac{1}{t\epsilon_{t}}\right).$
Therefore the best rate is obtained for $\eta=0.5+\delta/3$ and $\epsilon=0.75+\delta$:
$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(t^{-0.25+\delta}\right),$
where $\delta$ is arbitrarily small.
The above theorem holds for all $t$ if we choose $T_{1}$ and $T_{2}$ to be sufficiently large and then $\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq C\left(\eta_{t}+\frac{1}{t\epsilon_{t}}\right)$ for all $t$, for some $C>0$.
A smaller value of $\epsilon$ gives a better rate. The optimal rate is then dictated by the assumptions that $\eta_{t}$ is square-summable ($\eta>0.5$) and $\lim_{t\uparrow\infty}\epsilon_{t}^{2}/\eta_{t}^{3}=0$ ($2\epsilon>3\eta$).
@@ 4 Convergence Analysis
Let $\mathcal{G}(\boldsymbol{\alpha})$ be our $N$-player game with utility functions $u_{n}(\boldsymbol{x},\boldsymbol{\beta})$ and action set $\mathcal{X}$ where $\boldsymbol{\beta}=\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$. Consider a modified $N$-player game $\mathcal{G}^{\prime}$ with utility function $u^{\prime}_{n}(\boldsymbol{x})=r_{n}(\boldsymbol{x})$ for $n\in\{1,\ldots,N\}$ and action set $\{\boldsymbol{x}\mid\boldsymbol{x}\in\mathcal{X},\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}\}$. Under the assumption that there exists $\boldsymbol{x}$ such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$ and $h_{i}(\boldsymbol{x})<0$ for all $i=1,\ldots,M$, Theorem 1 along with equations (3.3) and (3.6) from [8] states that there exists a NE $\hat{\boldsymbol{x}}\in\mathcal{X}$ of $\mathcal{G}^{\prime}$, and $\boldsymbol{\mu}\in\mathbb{R}^{M},\boldsymbol{\lambda}\in\mathbb{R}^{K}$ such that
$F(\hat{\boldsymbol{x}})+\sum_{i=1}^{M}\nabla h(\hat{\boldsymbol{x}})\mu_{i}+\boldsymbol{A}^{\intercal}\boldsymbol{\lambda}=0$
@@
$F(\hat{\boldsymbol{x}},\boldsymbol{\alpha})+\sum_{i=1}^{M}\nabla h(\hat{\boldsymbol{x}})\mu_{i}=0$
@@
which implies that $\hat{\boldsymbol{x}}$ is a NE for the game $\mathcal{G}(\boldsymbol{\alpha})$, and hence $\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\hat{\boldsymbol{x}}$. We also know that $\boldsymbol{A}\hat{\boldsymbol{x}}=\boldsymbol{\ell^{*}}$. Hence there exists $\boldsymbol{\alpha}$ such that $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$, showing that $\mathcal{N}_{opt}$ is non-empty.
One of the main technical challenges in our analysis stems from the constrained actions represented by the convex and compact action sets. Indeed, with no constraints on the actions, and under certain assumptions on the matrix $\boldsymbol{A}$, we can show that $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ is strongly monotone instead of co-coercive (Lemma 4 (b)). This in turn shows that the mapping $g(\boldsymbol{\alpha})$ in Definition 2 is contractive. In this simplified case, both time-scales are fixed point iterations with contractive mappings and results from [11] can be applied. Therefore, the action sets call for a novel two time scale convergence analysis.
@@ 4.1 Non-expansive Mapping
where $\gamma=\min\{\mu/\|\boldsymbol{A}\|^{2},1\}$ is a constant chosen to satisfy $02\mu\gamma-\gamma^{2}\|\boldsymbol{A}\|^{2}>$ 0. Note that a fixed point of $g(\boldsymbol{\alpha})$ will satisfy our objective of $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$.
@@
Let $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ be the unique NE given $\boldsymbol{\alpha}\in\mathbb{R}^{K}$. Then,
@@
For a fixed $\boldsymbol{\alpha}$, $-F(\boldsymbol{x},\boldsymbol{\alpha})$ is strongly monotone on $\mathcal{X}$ and is Lipschitz continuous by Assumptions 1 and 2 respectively. Then [51, Theorem 2.3.3] states that for all $\boldsymbol{x}\in\mathcal{X}$, for some $L^{\prime}$
$\|\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha})\|\leq L^{\prime}\|\boldsymbol{x}-\Pi_{\mathcal{X}}(\boldsymbol{x}+F(\boldsymbol{x},\boldsymbol{\alpha}))\|.$
Hence for $\boldsymbol{\alpha}_{2},\boldsymbol{\alpha}_{1}\in\mathbb{R}^{K}$, we have
@@
$\leq L^{\prime}\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}))\|$
=$\stackrel{{\scriptstyle(a)}}{{=}}L^{\prime}\Big{\|}\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}))$
@@
where $L_{0}=L^{\prime}\|\boldsymbol{A}^{\intercal}\|$. Here equality (a) follows from [51, Proposition 1.5.8] which says
@@
Now let $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$ Since $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})$ is the NE corresponding to $\boldsymbol{\alpha}_{1}$, using [51, Proposition 1.4.2], we have for every $\boldsymbol{x}\in\mathcal{X}$ that $\left\langle\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\right\rangle\leq 0$. For $\boldsymbol{x}=\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})$
@@
$\leq\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}-(2\mu\gamma-\gamma^{2}\|\boldsymbol{A}\|^{2})\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\|\Pi_{\mathcal{X}}(\boldsymbol{x}_{t}+\eta_{t}(F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{M}_{t+1}))$
@@
Here equality (a) uses that $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\in\mathcal{X}$ and inequality (b) follows from since the projection operator $\Pi_{\mathcal{X}}(\cdot)$ for any convex set $\mathcal{X}$ is non-expansive.
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\eta_{t}^{2}\mathbb{E}[\|\boldsymbol{M}_{t+1}\|^{2}\mid\mathcal{F}_{t}]+\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
@@
Here equality (a) follows from the fact that $\alpha_{t+1}$ is $\mathcal{F}_{t}$ measurable and $\boldsymbol{M}_{t+1}$ is a Martingale difference sequence, hence making the term $\mathbb{E}[\langle\boldsymbol{M}_{t+1},\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\rangle\mid\mathcal{F}_{t}]=0$. Inequality (b) follows from Lemma 4 part (a), and (c) follows from $\boldsymbol{x}_{t}\in\mathcal{X}$ for some constant $c_{3}$.
@@
$\stackrel{{\scriptstyle}}{{\leq}}2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|\epsilon_{t}L_{0}\left\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\right\|$
$\;\;+2\eta_{t}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|\epsilon_{t}L_{0}\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|$
@@
$\leq(1-\frac{3}{2}\mu\eta_{t})\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]+\eta_{t}^{2}(2c_{1}^{2}+M_{c})$

### v3

@@ 2.1.1 Quadratic Global Objective
For a positive semi-definite matrix $\boldsymbol{G}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{\rho}\in\mathbb{R}^{Nd}$, the controller can set a global cost of the form $\Phi\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$. All the action profiles that are optima of $\Phi\left(\boldsymbol{x}\right)$ satisfy $\nabla\Phi\left(\boldsymbol{x}\right)=2\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}=0$. Hence, using our approach, we can set $\boldsymbol{A}=2\boldsymbol{G}$ and $\boldsymbol{\ell^{*}}=-\boldsymbol{\rho}$ to reach a global optimum.
@@ 2.1.2 Weighted Resource Allocation Games
In resource allocation, high congestion on a single resource increases the operation costs for the manager, can lead to system failures, and incurs suboptimal performance for the users. Therefore, load balancing is a key control objective for the manager who supervises the resource allocation protocol. In the electricity grid, the manager can be the utility company running Demand Side Management (DSM) to cut the production costs of energy at peak hours. In wireless networks, the manager can be an access point coordinating a protocol that minimizes the power consumption of the devices and the interference in the network. Other examples are load balancing in data-centers [53] and control of parking resources [54]. For distributed resource allocation, designing an efficient load balancing protocol becomes far more challenging.
@@
In wireless networks, the interference that the manager measures is a weighted sum of the transmission powers of the players, where the weights are the channel gains. In data centers, different users send jobs of different sizes to different servers (resources), so the load is a weighted sum of their actions (how many jobs to send where). A special case is resource allocation without weights, studied in [48].
@@ 3 Game Control Algorithm
Algorithm 1 Online Game Control
@@
[…]
@@
[…]
Our algorithm is a two time-scale stochastic approximation iteration. In the faster time-scale, dictated by the stepsize $\{\eta_{t}\}$, the agents update their action using gradient play, taking a step towards the NE corresponding to $\boldsymbol{\alpha}_{t}$. In the slower timescale, dictated by the stepsize $\{\epsilon_{t}\}$, the manager observes the constraint violation based on the players’ current action and updates $\boldsymbol{\alpha}_{t}$ to steer the player towards a desirable point.
Our players run SGD to optimize their rewards, and therefore are myopic and do not manipulate the manager using dynamic strategies. This behavior is expected from selfish players in a large-scale network or from cooperative players.
Selfish players are likely to maximize their rewards using a common algorithm such as gradient ascent [6]. In particular, when the players do not know their reward functions, small SGD steps allow for gradual learning, as opposed to best-response dynamics that can have large jumps. Nevertheless, our techniques can be applied to other distributed algorithms that are guaranteed to converge to NE.
The behavior of cooperative players is not something to model but to design [55]. The algorithm is used as a distributed protocol programmed into the devices in the network (e.g., WiFi). The challenge is then not to control the selfish behavior of the players, but to guide them towards a more efficient NE they do not have enough information to find otherwise (i.e., players do not know which $\boldsymbol{\alpha}$ would satisfy the linear constraints at NE). For this purpose, gradient play is an efficient distributed algorithm, that can lead to an efficient NE when combined with the manager’s algorithm in (8).
@@ 3.1 Performance Guarantees
The manager wants to converge to a suitable $\boldsymbol{\alpha}$ such that the linear constraints $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$ are satisfied at the NE corresponding to $\boldsymbol{\alpha}$, i.e., converge to the set
@@
We assume that Slater’s condition [56] for our problem holds, which allows us to prove that $\mathcal{N}_{opt}$ is non-empty:
@@
The set of action profiles is of the form $\mathcal{X}=\{\boldsymbol{x}\mid h_{i}(\boldsymbol{x})\leq 0,i=1,\ldots,M\}$ for some $M$ where each $h_{i}(\boldsymbol{x})$ is continuously differentiable and convex and there exists $\boldsymbol{x}\in\text{int}(\mathcal{X})$ (i.e., $h_{i}(\boldsymbol{x})<0$, $i=1,\ldots,M$) such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$.
@@
The stepsizes $\{\eta_{t}\}$ and $\{\epsilon_{t}\}$ are of the form $\eta_{t}=1/(t+T_{1})^{\eta}$ and $\epsilon_{t}=1/(t+T_{2})^{\epsilon}$ for some $T_{1},T_{2}>0$ and $0.5<\eta,\epsilon<1$ such that $\lim_{t\to\infty}\frac{\epsilon_{t}^{2}}{\eta_{t}^{3}}=0$.
The assumption that $0.5<\eta,\epsilon<1$ ensures that the stepsize sequences are monotonically decreasing and square-summable but not summable. The assumption that $\lim_{t\to\infty}\frac{\epsilon_{t}^{2}}{\eta_{t}^{3}}=0$ quantifies the required separation between the time-scales.
[…]
@@
Under Assumptions 1-5, $\boldsymbol{\alpha}_{t}$ converges to the set $\mathcal{N}_{opt}$, $\boldsymbol{x}_{t}$ converges to $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$, and with probability 1:
$\lim_{t\rightarrow\infty}\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}}.$
The next theorem gives the convergence rate for our algorithm. The bound holds for all $t>0$ if $T_{1},T_{2}>T_{0}$.
@@
Under Assumptions 1-5, there exists a $T_{0}>0$ and $C>0$ such that
$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq C\left(\eta_{t}+\frac{1}{t\epsilon_{t}}\right),$
for all $t>T_{0}$.
Remark 1.
A smaller value of $\epsilon$ gives a better rate, but it is constrained by the assumptions that $\eta_{t}$ is square-summable ($\eta>0.5$) and $\lim_{t\rightarrow\infty}\frac{\epsilon_{t}^{2}}{\eta_{t}^{3}}=0$ ($2\epsilon>3\eta$). Let $\delta>0$. Then for $\eta=0.5+\frac{\delta}{3}$ and $\epsilon=0.75+\delta$ we obtain
$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(t^{-0.25+\delta}\right).$
where $\delta>0$ can be chosen to be arbitrarily small.
@@ 4 Convergence Analysis
Let $\mathcal{G}(\boldsymbol{\alpha})$ be our game with utility functions $u_{n}(\boldsymbol{x},\boldsymbol{\beta})$ and set of action profiles $\mathcal{X}$ where $\boldsymbol{\beta}=\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$. Consider a modified game $\mathcal{G}^{\prime}$ with utility functions $u^{\prime}_{n}(\boldsymbol{x})=r_{n}(\boldsymbol{x})$ and set of action profiles $\{\boldsymbol{x}\mid\boldsymbol{x}\in\mathcal{X},\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}\}$. Under the assumption that there exists $\boldsymbol{x}$ such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$ and $h_{i}(\boldsymbol{x})<0$ for all $i=1,\ldots,M$, Theorem 1 from [4] states that there exists a NE $\hat{\boldsymbol{x}}\in\mathcal{X}$ of $\mathcal{G}^{\prime}$. Further, the Karush-Kuhn-Tucker conditions (equations (3.3) and (3.6) from [4]) state that there exist $\boldsymbol{\mu}\in\mathbb{R}^{M},\boldsymbol{\lambda}\in\mathbb{R}^{K}$ such that
$F(\hat{\boldsymbol{x}})+\sum_{i=1}^{M}\nabla h_{i}(\hat{\boldsymbol{x}})\mu_{i}+\boldsymbol{A}^{\intercal}\boldsymbol{\lambda}=0$
@@
$F(\hat{\boldsymbol{x}},\boldsymbol{\alpha})+\sum_{i=1}^{M}\nabla h_{i}(\hat{\boldsymbol{x}})\mu_{i}=0$
@@
[…]
A main technical challenge in our analysis is the convex and compact action sets. Indeed, with no constraints on the actions, and under certain assumptions on the matrix $\boldsymbol{A}$, we can show that $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ is strongly monotone instead of co-coercive (Lemma 4 (b)). This in turn shows that the mapping $g(\boldsymbol{\alpha})$ in Definition 2 is contractive. In this simplified case, both time-scales are fixed point iterations with contractive mappings and results from [7] can be applied. Therefore, the action sets call for a novel two time scale convergence analysis.
@@ 4.1 Non-expansive Mapping
where $\gamma=\min\{\mu/\|\boldsymbol{A}\|^{2},1\}$ is a constant chosen to satisfy $2\mu\gamma-\gamma^{2}\|\boldsymbol{A}\|^{2}>0$. Note that a fixed point of $g(\boldsymbol{\alpha})$ will satisfy our objective of $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$.
@@
Let $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ be the NE given $\boldsymbol{\alpha}\in\mathbb{R}^{K}$. Then,
@@
For a fixed $\boldsymbol{\alpha}$, $F(\boldsymbol{x},\boldsymbol{\alpha})$ is strongly monotone on $\mathcal{X}$ and is Lipschitz continuous by Assumptions 1 and 2 respectively. Then we state two properties of the NE $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$. First, [49, Proposition 1.4.2] states that the NE $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ corresponding to $\boldsymbol{\alpha}$ satisfies the variational inequality
$\left\langle\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}),\boldsymbol{\alpha}\right)\right\rangle\leq 0$ (11)
for all $\boldsymbol{x}\in\mathcal{X}$. Second, for a strongly monotone function, [49, Theorem 2.3.3] states that for all $\boldsymbol{x}\in\mathcal{X}$, for some $L^{\prime}$
$\|\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha})\|\leq L^{\prime}\|\boldsymbol{x}-\Pi_{\mathcal{X}}(\boldsymbol{x}+F(\boldsymbol{x},\boldsymbol{\alpha}))\|.$ (12)
Hence for $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$, we have
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}L^{\prime}\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}))\|$
$\stackrel{{\scriptstyle(b)}}{{=}}L^{\prime}\Big\|\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}))$
@@
where $L_{0}=L^{\prime}\|\boldsymbol{A}^{\intercal}\|$. Here inequality (a) follows from (12) and equality (b) follows from [49, Proposition 1.5.8] which states that the solution of the VI (11) satisfies
@@
Now let $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$. Then for $x=\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})$ in (11),
@@
$=\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}-(2\mu\gamma-\gamma^{2}\|\boldsymbol{A}\|^{2})\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\|\Pi_{\mathcal{X}}(\boldsymbol{x}_{t}+\eta_{t}(F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{M}_{t+1}))$
@@
Here equality (a) uses that $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\in\mathcal{X}$ and inequality (b) follows since the projection operator $\Pi_{\mathcal{X}}(\cdot)$ for any convex set $\mathcal{X}$ is non-expansive.
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\eta_{t}^{2}\mathbb{E}[\|\boldsymbol{M}_{t+1}\|^{2}\mid\mathcal{F}_{t}]+\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
@@
Here (a) uses $\mathbb{E}[\langle\boldsymbol{M}_{t+1},\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\rangle\mid\mathcal{F}_{t}]=0$, since $\alpha_{t+1}$ is $\mathcal{F}_{t}$ measurable and $\boldsymbol{M}_{t+1}$ is a Martingale difference sequence. Inequality (b) follows from Lemma 4 part (a), and (c) follows from $\boldsymbol{x}_{t}\in\mathcal{X}$ for some constant $c_{3}$.
@@
[…]
@@
$\leq\left(1-\frac{3}{2}\mu\eta_{t}\right)\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]+\eta_{t}^{2}(2c_{1}^{2}+M_{c})$